I need help with the following: 4x-9y = 61
10x+3y = 25
I am writing for my daughter. She has an IEP for math and is struggling with her work. She will write to you from this point forward, however, when you explain it would be helpful if you wrote your explanations as simply as possible. My daughter is also from another country. I am a music teacher and am no help whatsoever to her. Thanks, Mom (DB)

Answer by keksjr(7)

Using Elimination method:
4x-9y = 61 .......Eqtn I
10x+3y = 25.......Eqtn II
Multiply Eqtn I: by x10 to eliminate x: 40x-90y=610
Multiply Eqtn II: by x4 to eliminate x: 40x+12y=100
Subtract: 0-102y=510
Divide through by -102: y=-5
Substitute for y in any of the Eqtn above to find x:
Lets use Eqtn I: 4x-9y = 61
4x-9(-5) = 61
4x+45 = 61
4x = 61-45
4x = 16
Divide through 4:
x=4
(x,y)(4,-5)